# Exact calculation of degrees for lattice equations: a singularity approach

### v2

@@ 3. How to calculate degrees for lattice equations
Let us recall the definition of a prime factor of the relation $\Phi(B,C,D)=x^{*}$. Let
$\Phi(B,C,D)=\frac{\Phi_{1}(B,C,D)}{\Phi_{2}(B,C,D)}$
where $\Phi_{1},\Phi_{2}\in\mathbb{C}[B,C,D]$ are coprime. If $x^{*}\neq\infty$ (resp. $x^{*}=\infty$), a prime factor of the relation $\Phi(B,C,D)=x^{*}$ is a prime factor of the polynomial
$\Phi_{1}(B,C,D)-x^{*}\Phi_{2}(B,C,D)$
(resp. of the polynomial $\Phi_{2}(B,C,D)$).
It is well known that if $x^{*}\in\mathbb{P}^{1}(\mathbb{C})$ is generic, the polynomial $\Phi_{1}(B,C,D)-x^{*}\Phi_{2}(B,C,D)$ is irreducible since $\Phi_{1}$ and $\Phi_{2}$ are coprime. Thus, for any quad equation, there are at most finitely many $\mathbb{P}^{1}(\mathbb{C})$-values for which the equation does not satisfy the basic pattern condition.
@@
Suppose that a quad equation $\Phi$ satisfies the $\partial$-factor condition (Definition 3.6). Let $H\subset\mathbb{Z}^{2}$ be a domain and let $z$ be an arbitrary one of the initial variables. Let $z^{*}\in\mathbb{P}^{1}$ and consider the constant singularity pattern corresponding to $z=z^{*}$. Then, the set
@@
Suppose that a quad equation $\Phi$ satisfies the $\partial$-factor condition. Let $H\subset\mathbb{Z}^{2}$ be a domain and let $z$ be an arbitrary one of the initial variables. Let $z^{*}\in\mathbb{P}^{1}$ and consider the constant singularity pattern corresponding to $z=z^{*}$. Let $x_{t,n}(z^{*})=x^{*}\in\mathbb{P}^{1}(\mathbb{C})$ be the first constant singularity of this pattern and let $r$ be the multiplicity of $x_{t,n}(z^{*})=x^{*}$. Suppose that the equation satisfies the basic pattern condition for $x^{*}$ (Definition 3.7). Then, in the region $H_{\geq(t,n)}$, the constant singularity pattern corresponding to $z=z^{*}$ coincides with the $(t,n)$-translation of the basic pattern corresponding to $x^{\text{basic}}_{0,0}=x^{*}$ with all the multiplicities multiplied by $r$. That is, if $x^{\text{basic}}_{s,m}(x^{*})=\alpha$ is a constant singularity of multiplicity $\ell$ on $H^{\text{basic}}$, then $x_{t+s,n+m}(z^{*})=\alpha$ is a constant singularity of multiplicity $r\ell$ on $H$.
@@
Suppose that a quad equation $\Phi$ satisfies the $\partial$-factor condition. Let $H\subset\mathbb{Z}^{2}$ be a domain and let $z$ be an arbitrary one of the initial variables.
  • Let $S=\{\alpha_{1},\ldots,\alpha_{J}\}\subset\mathbb{P}^{1}(\mathbb{C})$, $S\neq\emptyset$. We focus on these values when calculating degrees. We call $\alpha_{1},\ldots,\alpha_{J}$ singular values.
  • Assume that we know all the singularity patterns that contain a singular value. Let $\beta_{1},\ldots,\beta_{I}\in\mathbb{P}^{1}(\mathbb{C})$ (resp. $\gamma_{1},\ldots,\gamma_{L}\in\mathbb{P}^{1}(\mathbb{C})$) be the starting values of the movable (resp. fixed) patterns we consider here.
@@
    be the multiplicity of $x^{\text{basic}}_{s,m}(\beta_{i})=\alpha_{j}$ in this pattern. The multiplicity $\operatorname{mult}^{\text{basic},\beta_{i}}_{s,m}(\alpha_{j})$ is considered to be 00 if $x^{\text{basic}}_{s,m}(\beta_{i})\neq\alpha_{j}$ or $(s,m)\notin H^{\text{basic}}$.
@@
    be the multiplicity of $x_{t,n}(\gamma_{\ell})=\alpha_{j}$ in this pattern. The multiplicity is considered to be 00 if $x_{t,n}(\gamma_{\ell})\neq\alpha_{j}$.
@@
Although $Z_{s,m}(\beta_{i})$ is used in Theorems 3.12 and 3.15, its definition is not given in this section. To calculate degrees with our theorems, however, the definition of $Z_{s,m}(\beta_{i})$ is not relevant. What is important here is the fact that $Z_{s,m}(\beta_{i})$ is a nonnegative integer defined in some rigorous way.
@@
If $t-s<0$ or $n-m<0$, then $\operatorname{mult}^{\text{basic},\beta_{i}}_{t-s,n-m}(\alpha_{j})$ is always 00. Therefore, the $\sum_{(s,m)\in H}$ in the degree relation can be replaced with $\sum_{(s,m)\in H_{\leq(t,n)}}$, which is always a finite sum.
@@
Let us use the same assumptions and notations as in Theorem 3.12.
@@
    If $I\geq J$, i.e., the number of singular values is equal to or greater than that of starting values, then it is impossible to completely compute degrees only from the degree relations.
@@
    If $\{\beta_{1},\ldots,\beta_{I}\}\subset S$ and $I<J$, then the degree relations completely determine all the degrees. That is, by solving a system of linear equations, one can compute degrees only from the degree relations.
@@
If the starting value of a movable pattern is $\beta$, then the pattern obviously contains $\beta$ as a value. Therefore, $\alpha\in S\setminus\{\beta_{1},\ldots,\beta_{I}\}$ implies that the equation does not have a movable pattern starting with $\alpha$. In short, to calculate degrees in our method, an equation needs to have a $\mathbb{P}^{1}(\mathbb{C})$-value that does not appear as the starting value of any movable pattern.
@@
Suppose that $\Phi$ satisfies the $\partial$-factor condition. Let $(t,n)\in H$ and suppose that $z^{*}\in\mathbb{P}^{1}$ generates a constant singularity at $(t,n)$. If $(s,m)\in H$ satisfies $s>t$ and
@@
all depends on the initial variable $w=x_{s,m-m^{\prime}}$ where $m^{\prime}$ is the least nonnegative integer satisfying $(s,m-m^{\prime})\in H_{0}$. In particular, they are not constant singularities, either (see Figure 5).
@@
[…]
Figure 5. Situation in Proposition 3.17. If the points marked with $\boxplus$” are not constant singularities, then those marked with “$+$” depend on $w$ and are not constant singularities, either. This procedure can easily be repeated and thus none of the points east to but not higher than the $\boxplus$-wall is a constant singularity. To prove this proposition, it is essential that the points marked with “$+$” lie outside the past light cone emanating from $(t,n)$.
@@
Suppose that $\Phi$ satisfies the $\partial$-factor condition. Let $x_{t,n}(z^{*})$ be the first constant singularity of a singularity pattern $z=z^{*}$ and let $(s,m)\in H_{\geq(t,n)}$. If
@@
Figure 6. Situation in Proposition 3.18. The point marked with “$\times$” is the first constant singularity. If none of the points marked with $\boxplus$” takes a constant singularity, then all the constant singularities of this pattern lie in the boxed region. If we consider a basic pattern, the assumption of the proposition coincides with what we check when we perform singularity confinement as in Example 2.1.
@@
Suppose that the equation satisfies the $\partial$-factor condition and let $x^{*}\in\mathbb{P}^{1}(\mathbb{C})$. Suppose that none of
@@
belongs to $\mathbb{P}^{1}(\mathbb{C})$. Then, the fixed pattern corresponding to $z=x^{*}$ is solitary. In particular, if the equation satisfies the basic pattern condition for $x^{*}$, then all the movable patterns starting with $x^{*}$ are solitary.
@@
Given a quad equation $\Phi$, one can find all the candidates for such $x^{*}$ by computing derivatives of $\Phi$ (and of $\frac{1}{\Phi}$ if necessary).
@@
The following proposition determines whether each candidate value obtained by Propositions  3.19 and 3.21 appears as the starting value of some movable pattern.
@@
has a prime factor other than
@@ 4. Proofs
Let $(t,n)\in H$ and suppose that $(t-1,n)$ and $(t,n-1)$ both belong to $H$. Let $x_{t,n}(z^{*})=x^{*}\in\mathbb{P}^{1}(\mathbb{C})$ be a constant singularity of multiplicity $r\in\mathbb{Z}_{>0}$ and suppose that $x_{t^{\prime},n^{\prime}}(z^{*})$ is not a constant singularity for any $(t^{\prime},n^{\prime})\in H_{\leq(t,n)}\setminus\{(t,n)\}$. If
@@
Let $(s,m)\in H^{\text{basic}}\setminus H^{\text{basic}}_{0}$ and let $\ell$ be the multiplicity of the constant singularity $x^{\text{basic}}_{s,m}(x^{*})=\alpha$ of the basic pattern (if it is not a constant singularity, we think of $\ell$ as 00). We show that the multiplicity of the constant singularity $x_{t+s,n+m}(z^{*})=\alpha$ is $r\ell$. We omit the proof for the case $x^{*}=\infty$ or $\alpha=\infty$, since in this case only the notation becomes complicated.
@@
Let $\epsilon$ be an infinitesimal parameter. Since $x^{\text{basic}}_{s,m}(x^{*})=\alpha$ is a constant singularity of multiplicity $\ell$, the expansion of $F(x^{*}+\epsilon;Y_{1},\ldots,Y_{N})$ is
$F(x^{*}+\epsilon;Y_{1},\ldots,Y_{N})=F_{\ell}(Y_{1},\ldots,Y_{N})\epsilon^{\ell}+O\left(\epsilon^{\ell+1}\right),$
@@
+++++++++++++++++++++$x^{*}$$z^{*}$
Figure 7. Situation in Lemma 4.1. The points marked with “$+$” can be thought of as if they were initial variables for a basic pattern since those points are algebraically independent over $\mathbb{C}$.
@@
Let $z^{*}\in\mathbb{P}^{1}$ and let $(t,n)\in H\setminus H_{0}$. If
@@
For $(t,n)\in H_{\geq(t_{0},n_{0})}$, we define a subset $\Delta_{t,n}\subset\mathbb{P}^{1}$ as
@@
The $\mathbb{P}^{1}(\mathbb{C})$ on the right-hand side is contained in the $\bigcup$-part unless it is the empty set, which happens only when $(t,n)=(t_{0},n_{0})$.
@@
It is clear by definition that $\mathbb{P}^{1}$ is covered as the direct sum of $\mathbb{P}^{1}(\mathbb{C})$ and $\Delta_{t,n}$:
$\mathbb{P}^{1}=\mathbb{P}^{1}(\mathbb{C})\sqcup\bigsqcup_{(t,n)\in H}\Delta_{t,n}.$
It will be revealed in Proposition 4.13 that this decomposition corresponds to where a constant singularity pattern starts.
@@
then $(t,n)\leq(s,m)$. In particular, if $x_{s,m}(z)$ takes a constant singularity at $z=z^{*}\in\Delta_{t,n}$, then $(t,n)\leq(s,m)$.
@@
In particular, if $x_{t,n}(z)$ takes a constant singularity at $z=z^{*}\in\Delta_{t,n}$, then the above transcendental degree is exactly $2$.
@@
Let us consider the remaining case, i.e., $t>t_{0}$. It is sufficient to show that $z^{*}$ is transcendental over $\mathbb{K}_{t-1,n}$. Let $n^{\prime}$ be the least positive integer that satisfies $(t,n-n^{\prime})\in H_{0}$ and let $w=x_{t,n-n^{\prime}}$. It follows from $t>t_{0}$ that $w\neq z$. Since $z^{*}\in\Delta_{t,n}$, $z^{*}$ depends on the initial variable $w$. Hence, $z^{*}$ is transcendental over $\mathbb{K}_{t-1,n}$, which completes the proof. ∎
@@
cannot be zero in the form $\frac{1}{\infty}$. Therefore, the first factor becomes 00 only if its numerator vanishes. Let $F=F(B,C,D)\in\mathbb{C}(B,C,D)$ be an arbitrary prime factor of the numerator of the rational function $\partial_{C}\Phi(B,C,D)$. By the $\partial$-factor condition, $F$ satisfies $F=B-\alpha$, $F=D-\alpha$ (for some $\alpha\in\mathbb{C}$) or $\partial_{C}F\neq 0$. If $F=B-\alpha$ or $F=D-\alpha$, then we have
@@
Since $F$ is an arbitrary prime factor of the numerator, we have
@@
If $x_{t,n}(z^{*})$ is the first singularity of a movable pattern, then $z^{*}$ belongs to $\Delta_{t,n}$. That is, the partition of $\mathbb{P}^{1}$ in Remark 4.7 corresponds to where a constant singularity pattern starts.
@@
Next, consider the case where $n=n_{0}$ or $t=t_{0}$. Since $(t,n)\neq(t_{0},n_{0})$, we may assume that $t=t_{0}$ and $n>n_{0}$. Note that $(t,n-1)$ belongs to $H_{\geq(t_{0},n_{0})}\cap H_{\leq(t,n)}$ but $(t-1,n)$ does not. Therefore, $\Delta_{t,n}$ in this case is written as
@@
Let $z^{*}\in\mathbb{P}^{1}$ and suppose that $x_{t,n}(z^{*})=x^{*}\in\mathbb{P}^{1}(\mathbb{C})$ is the first singularity of a movable pattern. Suppose that the equation satisfies the basic pattern condition for $x^{*}$. Then, $x_{t-1,n-1}(z^{*})$ is algebraic over $\mathbb{C}(x_{t-1,n}(z^{*}),x_{t,n-1}(z^{*}))$. In particular, $x_{t-1,n}(z^{*})$ and $x_{t,n-1}(z^{*})$ are algebraically independent over $\mathbb{C}$.
@@
there exists a prime factor $F=F(B,C,D)\in\mathbb{C}[B,C,D]$ of the relation $\Phi(B,C,D)=x^{*}$ such that
@@
For $(s,m)\in H$ and $\beta\in\mathbb{P}^{1}(\mathbb{C})$, we define a divisor $D_{s,m}(\beta)\in\operatorname{Div}\mathbb{P}^{1}$ as
@@
$x_{s,m}\colon\mathbb{P}^{1}\to\mathbb{P}^{1};\ z\mapsto x_{s,m}(z)$
@@
For $\gamma\in\mathbb{P}^{1}(\mathbb{C})$, we define a prime divisor $D(\gamma)\in\operatorname{Div}\mathbb{P}^{1}$ as

### v3

@@ 3. How to calculate degrees for lattice equations
If the polynomial $\Phi_{1}(B,C,D)-x^{*}\Phi_{2}(B,C,D)$ is irreducible, then $x^{*}$ satisfies the basic pattern condition. In many practical cases, this irreducibility holds for generic $x^{*}\in\mathbb{C}$ since $\Phi_{1}$ and $\Phi_{2}$ are coprime. For instance, if $\deg_{D}\Phi(B,C,D)=1$ (equivalent to the usual reversibility condition in discrete integrable systems, i.e., that the equation can be solved in the opposite direction by a rational function), then $\Phi_{1}(B,C,D)-x^{*}\Phi_{2}(B,C,D)$ is irreducible for generic $x^{*}$, as it is of degree $1$ in $D$. Similarly, if $\deg_{B}\Phi(B,C,D)=1$ or $\deg_{C}\Phi(B,C,D)=1$, the polynomial is irreducible for generic $x^{*}\in\mathbb{C}$.
However, there are some cases where $\Phi_{1}(B,C,D)-x^{*}\Phi_{2}(B,C,D)$ is reducible for generic $x^{*}$.
The first case occurs when
$\operatorname{trdeg}_{\mathbb{C}}\mathbb{C}\left(\Phi_{1}(B,C,D),\Phi_{2}(B,C,D)\right)=1,$
with $\operatorname{trdeg}_{K_{1}}K_{2}$ denoting the transcendental degree of a field extension $K_{1}\subset K_{2}$. In this situation, there exists a polynomial $Y=Y(B,C,D)$ such that $\Phi_{1}$ and $\Phi_{2}$ can be regarded as polynomials in $Y$ only: $\Phi_{1}=\Phi_{1}(Y)$, $\Phi_{2}=\Phi_{2}(Y)$. If either $\Phi_{1}(Y)$ or $\Phi_{2}(Y)$ has degree greater than $1$, then the polynomial $\Phi_{1}(Y)-x^{*}\Phi_{2}(Y)$ factorizes for generic $x^{*}$. However, any irreducible factorization must take the following form:
$\Phi_{1}(Y)-x^{*}\Phi_{2}(Y)=a(Y-b_{1})\cdots(Y-b_{m}).$
Since $\Phi$ depends on each of $B$, $C$, and $D$, so does $Y$. Therefore, the basic pattern condition still holds for generic $x^{*}$ in this case.
The second case occurs when there exist $\Psi_{1},\Psi_{2}\in\mathbb{C}[B,C,D]$ and $p\in\mathbb{Z}_{\geq 2}$ such that
$\Phi_{1}=\Psi^{p}_{1},\quad\Phi_{2}=\Psi^{p}_{2}.$
In this situation, the polynomial $\Phi_{1}(B,C,D)-x^{*}\Phi_{2}(B,C,D)$ trivially factorizes as
$\Phi_{1}-x^{*}\Phi_{2}=\Psi^{p}_{1}-x^{*}\Psi^{p}_{2}=\prod^{p-1}_{j=0}\left(\Psi_{1}-\zeta^{j}_{p}\sqrt[p]{x^{*}}\Psi_{2}\right),$
where $\zeta_{p}$ denotes a primitive $p$-th root of unity. Although each factor may further decompose, the analysis reduces to the case where $\Phi_{1}$ and $\Phi_{2}$ are replaced by $\Psi_{1}$ and $\Psi_{2}$, i.e., after removing the $p$-th power. Therefore, the basic pattern condition can be verified by considering the situation without such a $p$-th power structure.
These observations lead us to the following conjecture.
Conjecture 3.13.
For any quad equation, there exist at most finitely many values in $\mathbb{P}^{1}(\mathbb{C})$ for which the basic pattern condition fails.
@@
Suppose that a quad equation $\Phi$ satisfies the $\partial$-factor condition (Definition 3.9). Let $H\subset\mathbb{Z}^{2}$ be a domain and let $z$ be one of the initial variables. Let $z^{*}\in\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)$ and consider the constant singularity pattern corresponding to $z=z^{*}$. Then, the set
@@
Each basic pattern has only one starting point. A pattern with two or more starting points, such as shown in Figures 1 or 3, is undesirable in our study because it does not correspond to a basic pattern.
@@
Suppose that a quad equation $\Phi$ satisfies the $\partial$-factor condition. Let $H\subset\mathbb{Z}^{2}$ be a domain and let $z$ be one of the initial variables. Let $z^{*}\in\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)$ and consider the constant singularity pattern corresponding to $z=z^{*}$. Let $x_{t,n}(z^{*})=x^{*}\in\mathbb{P}^{1}(\mathbb{C})$ be the first constant singularity of this pattern and let $r$ be the multiplicity of $x_{t,n}(z^{*})=x^{*}$. Suppose that the equation satisfies the basic pattern condition for $x^{*}$ (Definition 3.11). Then, in the region $H_{\geq(t,n)}$, the constant singularity pattern corresponding to $z=z^{*}$ coincides with the $(t,n)$-translation of the basic pattern corresponding to $x^{\text{basic}}_{0,0}=x^{*}$ with all the multiplicities multiplied by $r$. That is, if $x^{\text{basic}}_{s,m}(x^{*})=\alpha$ is a constant singularity of multiplicity $\ell$ on $H^{\text{basic}}$, then $x_{t+s,n+m}(z^{*})=\alpha$ is a constant singularity of multiplicity $r\ell$ on $H$.
@@
Suppose that a quad equation $\Phi$ satisfies the $\partial$-factor condition. Let $H\subset\mathbb{Z}^{2}$ be a domain and let $z$ be one of the initial variables.
  • Let $\{\alpha_{1},\ldots,\alpha_{J}\}\subset\mathbb{P}^{1}(\mathbb{C}), SS\neq\emptyset$. We focus on these singular values when calculating degrees.
  • Assume that we know all the singularity patterns that contain a singular value $\alpha_{j}$, and assume in addition that the set of starting values of these patterns is finite. Let $\beta_{1},\ldots,\beta_{I}\in\mathbb{P}^{1}(\mathbb{C})$ (resp. $\gamma_{1},\ldots,\gamma_{L}\in\mathbb{P}^{1}(\mathbb{C})$) be the starting values of the movable (resp. fixed) patterns we consider here.
@@
    be the multiplicity of $x^{\text{basic}}_{s,m}(\beta_{i})=\alpha_{j}$ in this pattern. The multiplicity $\operatorname{mult}^{\text{basic},\beta_{i}}_{s,m}(\alpha_{j})$ is considered to be 0 if $x^{\text{basic}}_{s,m}(\beta_{i})\neq\alpha_{j}$ or $(s,m)\notin H^{\text{basic}}$.
@@
    be the multiplicity of $x_{t,n}(\gamma_{\ell})=\alpha_{j}$ in this pattern. The multiplicity is considered to be 0 if $x_{t,n}(\gamma_{\ell})\neq\alpha_{j}$.
@@
Although $Z_{s,m}(\beta_{i})$ is used in Theorems 3.17 and 3.20, its definition is not given in this section. To calculate degrees with our theorems, however, the definition of $Z_{s,m}(\beta_{i})$ is not relevant. What is important here is the fact that, for each $s,m,i$, $Z_{s,m}(\beta_{i})$ is a nonnegative integer defined in some rigorous way.
@@
If ts<0t-s<0 or nm<0n-m<0, then $\operatorname{mult}^{\text{basic},\beta_{i}}_{t-s,n-m}(\alpha_{j})$ is always 0. Therefore, the $\sum_{(s,m)\in H}$ in the degree relation can be replaced with $\sum_{(s,m)\in H_{\leq(t,n)}}$, which is always a finite sum.
@@
Under the same assumptions and notations as in Theorem 3.17, we have the following:
@@
    If $I\geq J$, i.e., the number of singular values is equal to or greater than that of starting values, then it is impossible to compute degrees only from the degree relations.
@@
    If $\{\beta_{1},\ldots,\beta_{I}\}\subset S$ and $I<J$, then the degree relations determine all the degrees. That is, by solving a system of linear equations, one can compute degrees only from the degree relations.
@@
If the starting value of a movable pattern is $\beta$, then the pattern contains $\beta$ as a value. Therefore, $\alpha\in S\setminus\{\beta_{1},\ldots,\beta_{I}\}$ implies that the equation does not have a movable pattern starting with $\alpha$. In short, to calculate degrees in our method, an equation needs to have a $\mathbb{P}^{1}(\mathbb{C})$-value that does not appear as the starting value of any movable pattern.
@@
Suppose that $\Phi$ satisfies the $\partial$-factor condition. Let $(t,n)\in H$ and suppose that $z^{*}\in\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)$ generates a constant singularity at $(t,n)$. If $(s,m)\in H$ satisfies $s>t$ and
@@
all depend on the initial variable $w=x_{s,m-m^{\prime}}$ where $m^{\prime}$ is the least nonnegative integer satisfying $(s,m-m^{\prime})\in H_{0}$. In particular, they are not constant singularities, either (see Figure 5).
@@
$x_{t,n}\boxplus\boxplus\boxplus\boxplus$$z^{*}$$w$
Figure 5. Situation in Proposition 3.22. If the points marked with $\boxplus$” are not constant singularities, then those marked with “++” depend on $w$ and are not constant singularities, either. This procedure can easily be repeated and thus none of the points east to but not higher than the $\boxplus$-wall is a constant singularity. To prove this proposition, it is essential that the points marked with “++” lie outside the past light cone emanating from $(t,n)$.
@@
Suppose that $\Phi$ satisfies the $\partial$-factor condition. Let $x_{t,n}(z^{*})$ be the first constant singularity of a singularity pattern and let $(s,m)\in H_{\geq(t,n)}$. If
@@
Figure 6. Situation in Proposition 3.23. The point marked with $\times$” is the first constant singularity. If none of the points marked with $\boxplus$” has a constant singularity, then all the constant singularities of this pattern lie in the boxed region. If we consider a basic pattern, the assumption of the proposition coincides with what we check when we perform singularity confinement as in Example 2.1.
@@
Suppose that the equation satisfies the $\partial$-factor condition and let $x^{*}\in\mathbb{P}^{1}(\mathbb{C})$. If none of
@@
belongs to $\mathbb{P}^{1}(\mathbb{C})$, then any singularity pattern starting from the value $x^{*}$ is solitary.
@@
Given a quad equation $\Phi$, one can find all $x^{*}$ that could generate non-solitary patterns by computing derivatives of $\Phi$ (and of $\frac{1}{\Phi}$ if necessary).
@@
The following proposition determines whether each candidate value obtained by Propositions  3.24 and 3.26 appears as the starting value of some movable pattern.
@@
has an irreducible factor other than
@@ 4. Proofs
Let $(t,n)\in H$ and suppose that (t1,n)(t-1,n) and (t,n1)(t,n-1) both belong to $H$. Let $x_{t,n}(z^{*})=x^{*}\in\mathbb{P}^{1}(\mathbb{C})$ be a constant singularity of multiplicity $r\in\mathbb{Z}_{>0}$ and suppose that $x_{t^{\prime},n^{\prime}}(z^{*})$ is not a constant singularity for any $(t^{\prime},n^{\prime})\in H_{\leq(t,n)}\setminus\{(t,n)\}$. If
@@
Let $(s,m)\in H^{\text{basic}}\setminus H^{\text{basic}}_{0}$ and let $\ell$ be the multiplicity of the constant singularity $x^{\text{basic}}_{s,m}(x^{*})=\alpha$ of the basic pattern (if it is not a constant singularity, we think of $\ell$ as 0). We show that the multiplicity of the constant singularity $x_{t+s,n+m}(z^{*})=\alpha$ is $r\ell$. We omit the proof for the case $x^{*}=\infty$ or $\alpha=\infty$, since in this case only the notation becomes complicated.
@@
Let $\varepsilon$ be an infinitesimal parameter. Since $x^{\text{basic}}_{s,m}(x^{*})=\alpha$ is a constant singularity of multiplicity $\ell$, the expansion of $F(x^{*}+\varepsilon;Y_{1},\ldots,Y_{N})$ is
$F(x^{*}+\varepsilon;Y_{1},\ldots,Y_{N})=F_{\ell}(Y_{1},\ldots,Y_{N})\varepsilon^{\ell}+O\left(\varepsilon^{\ell+1}\right),$
@@
$x^{*}$$z^{*}$
Figure 7. Situation in Lemma 4.1. The points marked with “++” can be thought of as if they were initial variables for a basic pattern since those points are algebraically independent over $\mathbb{C}$.
@@
Let $z^{*}\in\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)$ and let $(t,n)\in H\setminus H_{0}$. If
@@
For $(t,n)\in H_{\geq(t_{0},n_{0})}$, we define a subset $\Delta_{t,n}\subset\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)$ as
@@
The $\mathbb{P}^{1}(\mathbb{C})$ on the right-hand side is contained in the $\bigcup$-part unless that is the empty set, which happens only when $(t,n)=(t_{0},n_{0})$.
@@
It is clear by definition that $\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)$ is covered as the direct sum of $\mathbb{P}^{1}(\mathbb{C})$ and $\Delta_{t,n}$:
$\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)=\mathbb{P}^{1}(\mathbb{C})\sqcup\bigsqcup_{(t,n)\in H}\Delta_{t,n},$
[…]
@@
then $(t,n)\leq(s,m)$. In particular, if $x_{s,m}(z)$ has a constant singularity at $z=z^{*}\in\Delta_{t,n}$, then $(t,n)\leq(s,m)$.
@@
In particular, if $x_{t,n}(z)$ has a constant singularity at $z=z^{*}\in\Delta_{t,n}$, then the above transcendental degree is exactly $2$.
@@
Let us consider the remaining case, i.e., $t>t_{0}$. It is sufficient to show that $z^{*}$ is transcendental over $\mathbb{K}_{t-1,n}$. Let $n^{\prime}$ be the least positive integer that satisfies $(t,n-n^{\prime})\in H_{0}$ and let $w=x_{t,n-n^{\prime}}$. It follows from $t>t_{0}$ that $w\neq z$. Since $^{*}\in\Delta_{t,n}, zz^{*}$ depends on the initial variable $w$. Hence, $z^{*}$ is transcendental over $\mathbb{K}_{t-1,n}$, which completes the proof. ∎
@@
cannot be zero in the form $\frac{1}{\infty}$. Therefore, the first factor becomes 0 only if its numerator vanishes. Let $F=F(B,C,D)\in\mathbb{C}(B,C,D)$ be an arbitrary irreducible factor of the numerator of the rational function $\partial_{C}\Phi(B,C,D)$. By the $\partial$-factor condition, $F$ satisfies $F=B-\alpha$, $F=D-\alpha$ (for some $\alpha\in\mathbb{C}$) or $\partial_{C}F\neq 0$. If $F=B-\alpha$ or $F=D-\alpha$, then we have
@@
Since $F$ is an arbitrary irreducible factor of the numerator, we have
@@
If $x_{t,n}(z^{*})$ is the first singularity of a movable pattern, then $z^{*}$ belongs to $\Delta_{t,n}$. That is, the partition of $\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)$ in Remark 4.7 corresponds to where a constant singularity pattern starts.
@@
Next, consider the case where $n=n_{0}$ or $t=t_{0}$. Since $(t,n)\neq(t_{0},n_{0})$, we may assume that $t=t_{0}$ and $n>n_{0}$. Note that (t,n1)(t,n-1) belongs to $H_{\geq(t_{0},n_{0})}\cap H_{\leq(t,n)}$ but (t1,n)(t-1,n) does not. Therefore, $\Delta_{t,n}$ in this case is written as
@@
Definition 4.14 (Spontaneous Occurrence of $x_{t,n}(z^{*})=\alpha$).
For $z^{*}\in\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)$ and $\alpha\in\mathbb{P}^{1}$, we say that $x_{t,n}(z^{*})$ spontaneously becomes $\alpha$ if $x_{t,n}(z^{*})=\alpha$ and $z^{*}\in\Delta_{t,n}$. By Proposition 4.13, the above condition holds if and only if $x_{t,n}(z^{*})=\alpha$ is the first singularity of the pattern.
@@
Let $z^{*}\in\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)$ and suppose that $x_{t,n}(z^{*})=x^{*}\in\mathbb{P}^{1}(\mathbb{C})$ is the first singularity of a movable pattern. Suppose that the equation satisfies the basic pattern condition for $x^{*}$. Then, $x_{t-1,n-1}(z^{*})$ is algebraic over $\mathbb{C}(x_{t-1,n}(z^{*}),x_{t,n-1}(z^{*}))$. In particular, $x_{t-1,n}(z^{*})$ and $x_{t,n-1}(z^{*})$ are algebraically independent over $\mathbb{C}$.
@@
there exists an irreducible factor $F=F(B,C,D)\in\mathbb{C}[B,C,D]$ of the relation $\Phi(B,C,D)=x^{*}$ such that
@@
For $(s,m)\in H$ and $\beta\in\mathbb{P}^{1}(\mathbb{C})$, we define a divisor $D_{s,m}(\beta)\in\operatorname{Div}\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)$ as
@@
$x_{s,m}\colon\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)\to\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right);\ z\mapsto x_{s,m}(z)$
@@
For $\gamma\in\mathbb{P}^{1}(\mathbb{C})$, we define a prime divisor $D(\gamma)\in\operatorname{Div}\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)$ as